# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ III.1 Key PN term detection
Figure 4: Sensitivity comparisons between GR and BGR signals. Signal generation parameters are listed in Tables 2 and 1. The model demonstrates comparable performance across different false alarm rates (FAR).
@@
To quantify the model’s performance, we employ the sensitive distance metric Usman et al. (2016); Schäfer et al. (2023); Narola et al. (2023). As shown in Fig. 4, the model demonstrates comparable detection sensitivity for both GR and BGR signals across a range of FARs, validating its generalization capability. This result indicates that deep learning models trained on GR templates can reliably detect BGR signals, provided the PN deviations remain within realistic bounds.
@@ III.2 Generalization with extended PN deviations
Figure 5: Detection performance at a fixed luminosity distance of 600 Mpc for various signals. Chirp masses are below 30 $M_{\odot}$, with mass ratios between 0.9 and 1. GR signals show slightly better detection performance compared to BGR signals. Among BGR signals, the model generalizes best to $3\text{PN}^{(\ell)}$, while the All PN signals exhibit the lowest detection performance. However, overall detection abilities remain comparable across varying false alarm rates (FAR)
Figure 6: The network’s detection performance for different signals at varying luminosity distances, with one false alarm per year. The chirp quality of the signals are less than 30, with a quality ratio between 0.9 and 1. The network demonstrates the best recognition ability for GR signals, while its performance is the weakest for All PN signals. Overall, apart from a slight decline in performance for All PN signals, the network’s recognition ability across all signals shows minimal variation at different luminosity distances. Detection performance at varying luminosity distances with one false alarm per year. Chirp masses are below 30 $M_{\odot}$, with mass ratios between 0.9 and 1. The model shows the highest recognition ability for GR signals, with a slight decline in performance for All PN signals. However, for single PN terms, the model’s performance is largely consistent across all luminosity distances.
@@
Our analysis shows that even with a broad range of PN offsets, the model maintains robust performance across both GR and BGR signals. Notably, the most substantial performance differences occur when the chirp mass is below 30 $M_{\odot}$ and the mass ratio is between 0.9 and 1. As shown in Fig. 5, GR signals exhibit slightly better detection performance than BGR signals, with the model performing worst for the All PN term, which introduces the most significant deviations from GR. Nonetheless, the detection gap between GR and BGR signals is minimal, highlighting the impressive generalization capability of the deep learning model.
In Fig. 6, we extend the analysis to varying luminosity distances with one false alarm per year. As expected, GR signals achieve the highest recognition performance, but the model demonstrates consistent detection for BGR signals as well. The decline in performance for All PN signals remains minor, suggesting that even with large PN deviations, the model can generalize effectively.
@@ III.3 Validation with real gravitational-wave events
We analyze the background noise around GW150914, spanning September 12th to November 4th (GPS: 1126072320 to 1130680320), ensuring simultaneous operation of the Hanford and Livingston detectors. The effective observation time totals 21 days. As shown in Fig. 7, the metric for GW150914 is 6, marking it as the most significant event, with a threshold for one false alarm per year at 4.23. GW151012, initially a candidate event, also surpasses this threshold, indicating high confidence in its later identification as a gravitational wave signal.
@@ IV Conclusion
Next, we extended the PN range beyond current theoretical expectations to explore the limits of the model’s generalization. We introduced an “All PN” signal, incorporating all PN terms to simulate extreme deviations. The model maintained strong performance across various PN offsets and luminosity distances, with only a slight decrease in accuracy for the All PN case. This result underscores the model’s resilience, suggesting that it can reliably detect BGR signals even with substantial waveform deviations.
@@
In conclusion, this work establishes that deep learning models can not only detect gravitational wave signals with high generalization capability but also potentially identify the first BGR signals. This breakthrough opens new avenues for gravitational wave astronomy, with AI-driven methods at the forefront of discovering physics beyond GR.

### v2

@@ III.1 Key PN Term Detection
Figure 6: Sensitivity comparisons between GR and BGR signals. Signal generation parameters are listed in Tables 2 and 1.
@@
To quantify the model’s performance, we employ the sensitive distance metric Usman et al. (2016); Schäfer et al. (2023); Narola et al. (2023). The sensitive volume as a function of the false alarm rate (FAR) is defined by
[…]
where $\phi$ is the distribution of events over spatial coordinates $x$ and binary system parameters $\theta$, and $\epsilon$ is the detection efficiency of the pipeline at a false alarm rate $\mathcal{F}$ Schäfer et al. (2020). If the samples are drawn from within the redshifted volume $V_{0}$ Chen et al. (2021b), with
$V_{0}=\int_{z_{\min}}^{z_{\max}}\mathrm{d}z\,\frac{\mathrm{d}V_{c}}{\mathrm{d}z}\,\frac{1}{1+z}$ (4)
where $\mathrm{d}V_{c}/\mathrm{d}z$ is the differential comoving volume, then the sensitive volume is approximately
$V(\mathcal{F})\approx V_{0}\cdot\frac{N_{\text{draw}}}{N(\mathcal{F})}$ (5)
where $N(\mathcal{F})$ is the number of signals detected at a FAR less than $\mathcal{F}$ and $N_{\text{draw}}$ is the number of injected events. Finally, we calculate the radius of the sphere corresponding to the sensitive volume, thereby obtaining the sensitive distance.
As shown in Fig. 6, the model demonstrates comparable detection sensitivity for both GR and BGR signals across a range of FARs, validating its generalization capability. While the maximum sensitive distance ($\sim 1000$ Mpc at FAR = 10 year-1) is more conservative compared to some existing pipelines, this represents an inherent trade-off between maximum distance sensitivity and broader generalization capability—a deliberate choice in our approach to BGR signal detection. This result indicates that deep learning models trained on GR templates can reliably detect BGR signals, provided the PN deviations remain within realistic bounds—that is, parameter ranges where the waveforms remain mathematically stable and physically interpretable.
@@ III.2 Generalization with Extended PN Deviations
Figure 7: Detection performance at a fixed luminosity distance of 600 Mpc for various signals. Chirp masses are below 30 $M_{\odot}$, with mass ratios between 0.9 and 1. The SNR distribution of the signals is shown in Appendix C.
Figure 8: The network’s detection performance for different signals at varying luminosity distances, with one false alarm per year.
@@
Our analysis shows that even with a broad range of PN offsets, the model maintains robust performance across both GR and BGR signals. We focus our analysis on systems with chirp mass below 30 $M_{\odot}$ and mass ratios between 0.9 and 1, as this parameter space exhibits the most pronounced variations in detection performance compared to the broader mass range. As shown in Fig. 7, within this restricted parameter space, GR signals exhibit slightly better detection performance than BGR signals. Among the different PN modifications tested, the model generalizes best to $3\text{PN}^{(\ell)}$ corrections, showing the smallest performance gap relative to GR signals, while the “All PN” case—which incorporates cumulative effects from all PN orders—introduces the most significant deviations from GR and consequently exhibits the largest performance degradation. This hierarchy in generalization performance likely reflects how different PN corrections affect the waveform structure in ways that are more or less compatible with the features learned during training. Nonetheless, even in the worst case, the detection gap between GR and BGR signals remains minimal, highlighting the impressive generalization capability of the deep learning model.
In Fig. 8, we extend the analysis to varying luminosity distances with one false alarm per year. The chirp mass of the signals are less than $30M_{\odot}$, with a mass ratio between 0.9 and 1. The network demonstrates the best recognition ability for GR signals, while its performance is the weakest for All PN signals. Overall, apart from a slight decline in performance for All PN signals, the network’s recognition ability across all signals shows minimal variation at different luminosity distances. The model shows the highest recognition ability for GR signals, with a slight decline in performance for All PN signals. However, for single PN terms, the model’s performance is largely consistent across all luminosity distances. As expected, GR signals achieve the highest recognition performance, but the model demonstrates consistent detection for BGR signals as well. The decline in performance for All PN signals remains minor, suggesting that even with large PN deviations, the model can generalize effectively.
@@ III.3 Validation with Real Gravitational-Wave Events
We analyze the background noise around GW150914, spanning September 12th to November 4th (GPS: 1126072320 to 1130680320), ensuring simultaneous operation of the Hanford and Livingston detectors. The effective observation time totals 21 days. As shown in Fig. 9, the metric for GW150914 is 6, marking it as the most significant event, with a threshold for one false alarm per year at 4.23. GW151012, initially a candidate event, also surpasses this threshold, indicating high confidence in its later identification as a gravitational wave signal. While the overall number of triggers around GW150914 follows the background estimate, we note that the candidate associated with GW151012 exhibits a metric value significantly above the threshold (set for one false alarm per year). These observations may indicate either statistical fluctuations or potentially interesting candidates that warrant further investigation.
@@ IV Conclusion
Next, we extended the PN range beyond current theoretical expectations Abbott et al. (2019c) to explore the limits of the model’s generalization. We introduced an “All PN” signal, incorporating all PN terms to simulate extreme deviations. The model maintained strong performance across various PN offsets and luminosity distances, with only a slight decrease in accuracy for the All PN case. This result underscores the model’s resilience, suggesting that it can reliably detect BGR signals even with substantial waveform deviations.
@@
This study presents a proof-of-principle method whose strong suit is the generalization ability, but that requires further improvements to become competitive with specialized search pipelines. The promising results demonstrate that the fundamental approach has significant potential for future gravitational wave astronomy, particularly for detecting signals that might escape traditional template-based methods.
In conclusion, this work establishes that deep learning models can not only detect gravitational wave signals with high generalization capability but also potentially identify the first BGR signals. This improvement opens new avenues for gravitational wave astronomy, with AI-driven methods at the forefront of discovering physics beyond GR.
@@ Acknowledgements.
Appendix A Training and Validation Convergence
Figure 10: Training and validation loss plots with the use of separate ResNet models for H, L, and HL data streams.
As shown in Fig. 10, we have monitored both training and validation loss curves throughout the learning process, which confirm that the model converges properly without signs of overfitting. With respect to the use of separate ResNet models for the H, L, and HL data streams, our experiments show that training distinct models for each stream better captures the unique noise characteristics inherent to each detector.
Appendix B Model Outputs for GW151012 and GW151226
Figure 11: Network outputs for the detection of GW151012 and GW151226. The plots show the confidence metrics from our multi-channel analysis, demonstrating how different detection channels respond to these gravitational wave events. The horizontal axis represents time, and the vertical axis indicates the network’s confidence score for analysis.
In Fig. 11, we present example outputs for events such as those observed on GW151012 and GW151226, which helps clarify the decision-making process. All three channels successfully identified GW151012 with metrics exceeding our threshold, resulting in a high ranking statistic of 5.24, which strongly supports its classification as a gravitational wave event. For GW151226, while the two-channel network exceeded the threshold, the three-channel analysis yielded a lower-ranking statistic. This difference in detection patterns highlights the importance of our multi-channel approach: by incorporating multiple detection channels, we can better assess the confidence level of potential events and reduce the risk of missing genuine signals that might show varying strengths across different analysis channels.
Appendix C Signal SNR Distribution at Luminosity Distance of 600
Figure 12: Signal-to-noise ratio (SNR) distributions for different types of injected signals. The plot shows the network SNRs for GR and BGR events with various PN modifications, all generated with fixed luminosity distance (600 Mpc) and restricted chirp mass range ($<30M_{\odot}$).
In our evaluation of specific BGR events, we fix the luminosity distance at 600 Mpc and restrict the chirp masses to values below $30\,M_{\odot}$ with mass ratios between 0.9 and 1. This selection maintains a consistent signal-to-noise ratio (SNR) across events, as shown in Fig. 12 so that the impact of PN modifications can be directly compared. In our experimental setup, the uniform SNR ensures that any reduced detection performance is attributed primarily to waveform deviations.